OpenStudy (anonymous):

PLZ HELP#VERY CONFUSED rewrite each expression in term with no power greater than 1 cos^3 theta

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

\[\cos ^{3}\theta \]

OpenStudy (anonymous):

The only way I can think of doing this is using complex numbers. Using Euler's formula we know that:\[\cos \theta =\frac{e^{i\theta}+e^{-i\theta}}{2}\]Therefore:\[\cos^3\theta=\left(\frac{e^{i\theta}+e^{-i\theta}}{2}\right)^3\]\[=\frac{1}{2^3}\left(e^{3i\theta}+3e^{2i\theta}e^{-i\theta}+3e^{i\theta}e^{-2i\theta}+e^{-3i\theta}\right)\]\[=\frac{1}{2^3}\left(e^{3i\theta}+3e^{i\theta}+3e^{-i\theta}+e^{-3i\theta}\right)\]\[=\frac{1}{2^3}\left(e^{3i\theta}+e^{-3i\theta}\right)+\frac{3}{2^3}\left(e^{i\theta}+e^{-i\theta}\right)\]\[\frac{1}{4}\left(\frac{e^{3i\theta}+e^{-3i\theta}}{2}\right)+\frac{3}{4}\left(\frac{e^{i\theta}+e^{-i\theta}}{2}\right)\]Using Euler's formula again backwards yields:\[=\frac{1}{4}\cos 3\theta+\frac{3}{4}\cos \theta\]